# A Framework for Safe Probabilistic Invariance Verification of Stochastic Dynamical Systems

### v1

@@ 8.2 Semi-definite Programming Implementation
$\begin{split}&\textbf{SDP0}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,1}-\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]\in\sum[\bm{x}],\\ v(\bm{x})-1-s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{0}(\bm{x})\in\sum[\bm{x}],s_{1}(\bm{x})\in\sum[\bm{x}].\end{cases}\end{split}$
@@
$\begin{split}&\textbf{SDP1}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),s_{i}(\bm{x}),i=0,1}-\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+s_{2}(\bm{x})\widehat{h}(\bm{x})-s_{3}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})+s_{4}(\bm{x})\widehat{h}(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,4.\end{cases}\end{split}$
@@
$\begin{split}&\textbf{SDP2}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),\epsilon_{2},s_{i}(\bm{x}),i=0,\ldots,2}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}\epsilon_{2}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ -\delta-\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]+v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})+s_{2}(\bm{x})\widehat{h}(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{2}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,2,\end{cases}\end{split}$
@@
$\begin{split}&\textbf{SDP3}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),w(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,4}-\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]+w(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+s_{3}(\bm{x})\widehat{h}(\bm{x})-s_{4}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,4.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP4}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{2},v(\bm{x}),w(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,4}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})-1+\epsilon_{2}+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ \mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]-v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ -v(\bm{x})+\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ 1-v(\bm{x})+s_{3}(\bm{x})\widehat{h}(\bm{x})-s_{4}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{2}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,4.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP5}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,1,2}-\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ -\mathcal{A}v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+p(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,1,2.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP6}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),u(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,1,2}-\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ -\mathcal{A}v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathcal{A}u(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+p(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,1,2.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP7}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{2},v(\bm{x}),u(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,1,2}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})-1+\epsilon_{2}+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ \mathcal{A}v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \mathcal{A}u(\bm{x})-v(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ 1-v(\bm{x})+p(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{2}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,1,2.\end{cases}\end{split}$

### v2

@@ Appendix B Optimizations for Computing Lower and Upper Bounds
$\begin{split}&\textbf{Op0}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1},\qquad\quad~{}~{}\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))],~{}\forall\bm{x}\in\mathbb{R}^{n}\\ v(\bm{x})\geq 1,\qquad\quad~{}~{}\forall\bm{x}\in\mathbb{R}^{n}\setminus\mathcal{X}\\ v(\bm{x})\geq 0,\qquad\qquad\quad~{}~{}\forall\bm{x}\in\mathbb{R}^{n}\\ \epsilon_{1}\geq 0\end{cases}\end{split}$
@@
$\begin{split}&\textbf{Op1}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1},\qquad\quad~{}~{}\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))],~{}\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1,\qquad\quad~{}~{}~{}\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ v(\bm{x})\geq 0,\qquad\qquad\quad~{}~{}\forall\bm{x}\in\widehat{\mathcal{X}}\\ \epsilon_{1}\geq 0\end{cases}\end{split}$
@@
$\begin{split}&\textbf{Op2}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),\epsilon_{2}}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq\epsilon_{2},\qquad\quad~{}~{}\forall\bm{x}\in\mathcal{X}_{0}\\ \mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta,\\ \qquad\qquad\qquad\qquad~{}\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 0,\qquad\qquad\,\forall\bm{x}\in\widehat{\mathcal{X}}\\ \epsilon_{2}\geq 0\end{cases}\end{split}$
$\begin{split}&\textbf{Op3}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),w(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1},\qquad\qquad\qquad~{}~{}~{}\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))],\qquad\quad~{}~{}\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x}),\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1,\qquad\qquad\qquad\quad\,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ \epsilon_{1}\geq 0\end{cases}\end{split}$
$\begin{split}&\textbf{Op4}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),w(\bm{x}),\epsilon_{2}}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\geq 1-\epsilon_{2},\qquad\qquad\qquad\quad\,\forall\bm{x}\in\mathcal{X}_{0}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))],\qquad\qquad\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x}),~{}\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1,\qquad\qquad\qquad\quad~{}\,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\\ \epsilon_{2}\geq 0\end{cases}\end{split}$
$\begin{split}&\textbf{Op5}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1},&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0,&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 0,&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 1,&\forall\bm{x}\in\partial\mathcal{X}\\ \epsilon_{1}\geq 0\end{cases}\end{split}$
$\begin{split}&\textbf{Op6}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),u(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1},\quad\,\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0,\quad\quad~{}\,\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathcal{A}u(\bm{x}),\quad\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1,\quad\quad~{}~{}\forall\bm{x}\in\partial{\mathcal{X}}\\ \epsilon_{1}\geq 0\end{cases}\end{split}$
$\begin{split}&\textbf{Op7}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),u(\bm{x}),\epsilon_{2}}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\geq 1-\epsilon_{2},\quad\,\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\geq 0,\quad\quad~{}\,\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathcal{A}u(\bm{x}),\quad\forall\bm{x}\in\mathcal{X}&\\ v(\bm{x})\leq 1,\quad\quad~{}~{}\forall\bm{x}\in\partial{\mathcal{X}}\\ \epsilon_{2}\geq 0\end{cases}\end{split}$
Appendix C Semi-definite Programming Implementation
$\begin{split}&\textbf{SDP0}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,1}-\epsilon_{1}\\ &\text{s.t.~{}}\\ &\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]\in\sum[\bm{x}],\\ v(\bm{x})-1-s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{0}(\bm{x})\in\sum[\bm{x}],s_{1}(\bm{x})\in\sum[\bm{x}].\end{cases}\end{split}$
$\begin{split}&\textbf{SDP1}~{}~{}~{}~{}~{}~{}~{}\min_{\epsilon_{1},v(\bm{x}),s_{i}(\bm{x}),i=0,1}-\epsilon_{1}\\ &\text{s.t.~{}}\\ &\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+s_{2}(\bm{x})\widehat{h}(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-s_{3}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})+s_{4}(\bm{x})\widehat{h}(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,4.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP2}~{}\min_{v(\bm{x}),\epsilon_{2},s_{i}(\bm{x}),i=0,\ldots,2}\epsilon_{2}\\ &\text{s.t.~{}}\\ &\begin{cases}\epsilon_{2}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ -\delta-\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]+v(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})+s_{2}(\bm{x})\widehat{h}(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{2}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,2,\end{cases}\end{split}$
@@ Appendix C Semi-definite Programming Implementation
$\begin{split}&\textbf{SDP3}~{}\min_{\epsilon_{1},v(\bm{x}),w(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,4}-\epsilon_{1}\\ &\text{s.t.~{}}\\ &\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]+w(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+s_{3}(\bm{x})\widehat{h}(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-s_{4}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,4.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP4}~{}\min_{\epsilon_{2},v(\bm{x}),w(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,4}\epsilon_{2}\\ &\text{s.t.~{}}\\ &\begin{cases}v(\bm{x})-1+\epsilon_{2}+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ \mathbb{E}^{\infty}[v(\bm{\phi}^{\bm{x}}_{\pi}(1))]-v(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ -v(\bm{x})+\mathbb{E}^{\infty}[w(\bm{\phi}^{\bm{x}}_{\pi}(1))]-w(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ 1-v(\bm{x})+s_{3}(\bm{x})\widehat{h}(\bm{x})\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-s_{4}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{2}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,\ldots,4.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP5}~{}\min_{\epsilon_{1},v(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,1,2}-\epsilon_{1}\\ &\text{s.t.~{}}\\ &\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ -\mathcal{A}v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+p(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,1,2.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP6}~{}\min_{\epsilon_{1},v(\bm{x}),u(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,1,2}-\epsilon_{1}\\ &\text{s.t.~{}}\\ &\begin{cases}1-\epsilon_{1}-v(\bm{x})+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ -\mathcal{A}v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-\mathcal{A}u(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ v(\bm{x})-1+p(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{1}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,1,2.\end{cases}\end{split}$
$\begin{split}&\textbf{SDP7}~{}\min_{\epsilon_{2},v(\bm{x}),u(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,1,2}\epsilon_{2}\\ &\text{s.t.~{}}\\ &\begin{cases}v(\bm{x})-1+\epsilon_{2}+s_{0}(\bm{x})g(\bm{x})\in\sum[\bm{x}],\\ \mathcal{A}v(\bm{x})+s_{1}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \mathcal{A}u(\bm{x})-v(\bm{x})+s_{2}(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ 1-v(\bm{x})+p(\bm{x})h(\bm{x})\in\sum[\bm{x}],\\ \epsilon_{2}\geq 0,\\ s_{i}(\bm{x})\in\sum[\bm{x}],i=0,1,2.\end{cases}\end{split}$